# compat=1.1.0

### v1

@@ I Introduction
One intriguing extension of the freezeout idea is that, within the dark sector, CDM particles annihilate into heavier states, through a process that would be forbidden at zero temperature, but is allowed at finite temperature. This mechanism is known as forbidden dark matter (FDM) [14, 15]. The thermal suppression in the FDM mechanism allows for realistically large relic densities at freezeout, but with smaller CDM masses and a broad range of self-interaction strengths.
Here, we propose a natural realization of FDM within dEFT. This effective theory extends the conventional chiral Lagrangian with its pNGB fields, $\pi$, to include a dilaton field, $\chi$. It has been extensively studied, in our Refs. [16, 17, 18, 19, 20, 21] and in Refs. [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. See also the precursors in Refs. [40, 41] and Ref. [42]. dEFT emerges as the low-energy description of certain underlying gauge theories amenable to lattice studies. One notable example is the $SU(3)$ gauge theory with $N_{f}=8$ fundamental Dirac fermions [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Lattice studies of this theory have reported evidence for the presence of a light singlet, scalar particle, along with $N^{2}_{f}-1$ pNGBs, in the accessible range of fermion masses.
@@ II Dilaton Effective Field Theory
To compute the dark-matter relic density, we need to know the cross section for the forbidden process, $\pi\pi\rightarrow\chi\chi$, and its inverse, $\chi\chi\rightarrow\pi\pi$. These depend on the parameters of dEFT: the masses, $M_{d}^{2}$ and $M_{\pi}^{2}$, the decay constants, $F_{d}^{2}$ and $F_{\pi}^{2}$, the pNGB count, $N_{\pi}=N_{f}^{2}-1$, and the dimensionless parameters $y$ and $\gamma$.
Lattice studies of specific gauge theories, which provide ultraviolet completions for dEFT, yield (model-dependent) information about some of the dEFT parameters. For example, in the $\mathrm{SU}(3)$ gauge theory with $N_{f}=8$ fundamental fermions, fits of dEFT to lattice data constrain the parameter $y$ to be close to 2 [35, 36, 37]; for simplicity, we set $y=2$ in the following, although recovering the more general expressions is straightforward. The parameter $\gamma$ is less well determined, and we keep it general.
@@ III Relic Density
where we have neglected corrections of order $1/x=T/M_{\pi}$. The four Feynman diagrams that contribute are shown in Fig. 1. We note that $\gamma$ cannot be too large or else the cross section, Eq. (9), would exceed unitarity bounds, signaling that the dEFT is outside its domain of validity.
[…]
Figure 1: The Feynman diagrams that contribute to the forbidden annihilation of two pNGBs, $\pi$, into two dilatons, $\chi$, within dEFT at leading order.
@@
Figure 2: In red, the numerical solution to Eq. (12) for the yield function $Y_{\pi}$ is shown for the following choice of model parameters: $M_{\pi}=1\,\text{GeV}$, $M_{\pi}/F_{\pi}=4$, $F^{2}_{\pi}/F^{2}_{d}=0.1$, $\Delta=0.3$, $\gamma=3$ and $y=2$. The equilibrium yield $Y_{\pi}^{\text{eq}}$ is plotted as the solid black line for comparison. The freezeout temperature, $x_{f}$, is represented as the vertical dashed line. The yield at late times (large $x$) is plotted as the dashed horizontal line.
@@
The form of this constraint depends on the parameter sensitivity of $Y_{\pi}(\infty)$, determined through the numerical solution to Eq. (12). In Fig. 3, we plot the allowed values of $M_{\pi}$ versus $\Delta$ for three fixed values of the ratio $M_{\pi}/F_{\pi}$. These correspond to three fixed values of the effective expansion parameter of dEFT.
Fig. 3 can be understood qualitatively by inspection of Eqs. (12) and (13), which determine $Y_{\pi}(\infty)$. It can be argued that this quantity grows approximately linearly with $M_{\pi}$ and exponentially with $\Delta$. This behavior can be exhibited explicitly if Eq. (12) is solved analytically, which is possible if the evolution takes place at temperatures $T$ where $g(T)$ and $h(T)$ are approximately constant. These functions however vary rapidly for temperatures around the QCD confinement scale, giving rise to the kinks visible in Fig. 3 for $M_{\pi}\sim 3$ GeV.
@@
Using the numerical solution of Eq. (12), we can also determine the freezeout temperature, $T_{f}$. We take this to be the temperature below which the comoving number density of pNGBs, $Y_{\pi}(x)$, begins to rise significantly above the value it would have if it were in thermal equilibrium with the rest of the SM $Y_{\pi}^{\text{eq}}(x)$, indicating that the pNGBs have decoupled from the SM bath. Following the convention of [73], we take $Y_{\pi}(x_{f})=2.5Y^{\text{eq}}_{\pi}(x_{f})$, where $x_{f}=M_{\pi}/T_{f}$. See Figs. 2 and 4.
@@ III.1 Subdominant $3\rightarrow 2$ Processes
In addition to the processes of Fig. 1, the population of the dark-matter pNGBs can be depleted through $3\rightarrow 2$ processes, as in the composite SIMP models of Refs. [59, 60, 61, 62]. There, with the dark matter described only by pNGBs (no dilaton), the relevant five-point interaction arises from the Wess-Zumino-Witten term. In our framework, we must also account for the $3\rightarrow 2$ process, $\pi\pi\pi\rightarrow\chi\pi$, with an outgoing dilaton, which arises at leading order in dEFT. It is unsuppressed thermally, as long as $M_{d}<2M_{\pi}$.
@@ IV Interaction with the SM
Equation (24) leads to an upper bound on $\Gamma_{\chi\rightarrow SM}$ since the dominant contribution to the left-hand side comes from an s-channel dilaton exchange as depicted in Fig. 5. The general “resonance” expression for this cross section is
@@
[…]
Figure 5: pNGBs can annihilate directly into final states containing $n$-SM particles through interactions shown in Eq. (22).

### v4

@@ I Introduction
One intriguing extension of the freezeout idea is that, within the dark sector, CDM particles annihilate into heavier states, through a process that would be forbidden at zero temperature, but is allowed at finite temperature. This mechanism is known as forbidden dark matter (FDM). See Refs. [14, 15], and [16, 17] for applications. The thermal suppression in the FDM mechanism allows for realistically large relic densities at freezeout, but with smaller CDM masses and a broad range of self-interaction strengths.
Here, we propose a natural realization of FDM within dEFT. This effective theory extends the conventional chiral Lagrangian with its pNGB fields, $\pi$, to include a dilaton field, $\chi$. It has been extensively studied, in our Refs. [18, 19, 20, 21, 22, 23] and in Refs. [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. See also the precursors in Refs. [43, 44] and Ref. [45]. dEFT emerges as the low-energy description of certain underlying gauge theories amenable to lattice studies. One notable example is the $SU(3)$ gauge theory with $N_{f}=8$ fundamental Dirac fermions [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. Lattice studies of this theory have reported evidence for the presence of a light singlet, scalar particle, along with $N^{2}_{f}-1$ pNGBs, in the accessible range of fermion masses.
@@ II Dilaton Effective Field Theory
To compute the dark-matter relic density, we need to know the cross section for the forbidden process, $\pi\pi\rightarrow\chi\chi$, and its inverse, $\chi\chi\rightarrow\pi\pi$. These depend on the parameters of dEFT: the masses, $M_{d}^{2}$ and $M_{\pi}^{2}$, the decay constants, $F_{d}^{2}$ and $F_{\pi}^{2}$, the number of pNGBs, which we call $N_{\pi}$, and the dimensionless parameters $y$ and $\gamma$.
Lattice studies of specific gauge theories, which provide ultraviolet completions for dEFT, yield (model-dependent) information about some of the dEFT parameters. For example, in the $\mathrm{SU}(3)$ gauge theory with $N_{f}=8$ fundamental fermions, $N_{\pi}=N^{2}_{f}-1=63$ and fits of dEFT to lattice data constrain the parameter $y$ to be close to 2 [37, 38, 39], a value also supported by theoretical arguments [71]. For simplicity, we set $y=2$ in the following, although recovering the more general expressions is straightforward. The parameter $\gamma$ is less well determined, and we keep it general.
@@ III Relic Density
where we have neglected corrections of order $1/x=T/M_{\pi}$. We note that $\gamma$ cannot be too large or else the cross section, Eq. (9), would exceed unitarity bounds, signaling that the dEFT is outside its domain of validity.
@@
Figure 1: In red, the numerical solution to Eq. (12) for the yield function $Y_{\pi}$ is shown for the choice $M_{\pi}=1\,\text{GeV}$ with model parameters $M_{\pi}/F_{\pi}=4$, $F^{2}_{\pi}/F^{2}_{d}=0.1$, $\Delta=0.3$, $\gamma=3$ and $y=2$. The equilibrium yield $Y_{\pi}^{\text{eq}}$ is plotted as the solid black line for comparison. The freezeout temperature, $x_{f}$, is represented as the vertical dashed line. The yield at late times (large $x$) is plotted as the dashed horizontal line.
@@
The form of this constraint depends on the parameter sensitivity of $Y_{\pi}(\infty)$, determined through the numerical solution to Eq. (12). A range of possible values of the dark-matter mass scale $M_{\pi}$ emerges depending on the dEFT parameters. In Fig. 2, we plot the value of $M_{\pi}$ versus $\Delta$ for three values of the ratio $M_{\pi}/F_{\pi}$, corresponding to three values of the effective expansion parameter. For a wide range of dEFT parameters, we are led to a mass scale $M_{\pi}$ in the broad GeV range.
Figure 2 can be understood qualitatively by inspection of Eqs. (12) and (13), which determine $Y_{\pi}(\infty)$. It can be argued that this quantity grows approximately linearly with $M_{\pi}$ and exponentially with $\Delta$. This behavior can be exhibited explicitly if Eq. (12) is solved analytically, which is possible if the evolution takes place at temperatures $T$ where $g(T)$ and $h(T)$ are approximately constant. These functions however vary rapidly for temperatures around the QCD confinement scale, giving rise to the kinks visible in Fig. 2 for $M_{\pi}\sim 3$ GeV.
@@
Using the numerical solution of Eq. (12), we can also determine the freezeout temperature, $T_{f}$. We take this to be the temperature below which the comoving number density of pNGBs, $Y_{\pi}(x)$, begins to rise significantly above the value it would have if it were in thermal equilibrium with the rest of the SM $Y_{\pi}^{\text{eq}}(x)$, indicating that the pNGBs have decoupled from the SM bath. Following the convention of [77], we take $Y_{\pi}(x_{f})=2.5Y^{\text{eq}}_{\pi}(x_{f})$, where $x_{f}=M_{\pi}/T_{f}$. We plot $x_{f}$ as a vertical dashed line in Fig. 1.
In Fig. 3, we show the dependence of $x_{f}$ on $M_{\pi}$. The figure indicates this dependence is approximately linear when viewed using a log scale for $M_{\pi}$, which can be explained using Eq. (14). Since $\Omega_{\text{CDM}}h^{2}$ is being fixed to its observed value, it implies that $M_{\pi}Y_{\pi}(\infty)=0.4\,\text{eV}$. Furthermore, $Y_{\pi}(\infty)\sim n^{\text{eq}}_{\pi}(T_{f})/s(T_{f})\sim x_{f}^{3/2}e^{-x_{f}}$ up to constants. Inverting this relationship approximately yields $x_{f}\sim\text{const}+\log(M_{\pi}/\text{GeV})$. Note that $n^{\text{eq}}_{\pi}(T)$ carries no direct dependence on the ratio $M_{\pi}/F_{\pi}$, explaining the observed lack of variation with this quantity. The kinks visible around $M_{\pi}\approx 3$ GeV arise due to the abrupt change in relativistic degrees of freedom $h(T)$ around the QCD confinement scale.
@@ III.1 Subdominant $3\rightarrow 2$ Processes
In addition to the $2\rightarrow 2$ annihilations of dark-matter pNGBs to dilatons, the population of pNGBs can also be depleted through $3\rightarrow 2$ processes, as in the composite SIMP models of Refs. [62, 63, 64, 65]. There, with the dark matter described only by pNGBs (no dilaton), the relevant five-point interaction arises from the Wess-Zumino-Witten term. In our framework, we must also account for the $3\rightarrow 2$ process, $\pi\pi\pi\rightarrow\chi\pi$, with an outgoing dilaton, which arises at leading order in dEFT. It is unsuppressed thermally, as long as $M_{d}<2M_{\pi}$.
@@ IV Interaction with the SM
Equation (24) leads to an upper bound on $\Gamma_{\chi\rightarrow SM}$ since the dominant contribution to the left-hand side comes from an s-channel dilaton exchange. The general “resonance” expression for this cross section is
@@ Acknowledgements.
The work of MP has been supported in part by the STFC Consolidated Grants No. ST/P00055X/1, ST/T000813/1, and ST/X000648/1. MP received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 813942.